OpenStudy (anonymous):

Find by hand the intervals where the function is increasing and decreasing. Use this information to sketch a graph. y = x^4 - 8x^2 + 1

OpenStudy (anonymous):

@abb0t can u help?

OpenStudy (anonymous):

take the derivative and check the signs of that derivative

OpenStudy (anonymous):

\[y = x^4 - 8x^2 + 1 \] \[y'=4x^3-16x=4x(x^2-4)=x(x+2)(x-2)\] is the first step

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

this will be negative on \((-\infty, -2)\) and also on \((0,2)\) is that fairly clear?

OpenStudy (amistre64):

lol, find "by hand" .... what does that even mean?

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

if i have to sit here and start carving a function on a stone tablet .. im gonna get real upset :/

OpenStudy (anonymous):

good question i was using my brain

OpenStudy (anonymous):

in the old days they wrote in the sand

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so function is decreasing on ((-\infy, -2)\) and also on \((0,2)\) increasing elsewise

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

make that decreasing on \((-\infty, -2)\)

OpenStudy (anonymous):

Okay i see

OpenStudy (amistre64):

satelitte is cheating, i think you have to do this by first principles, not those modern derivative rules ...

OpenStudy (anonymous):

ohh lol okay